OpenStudy (anonymous):

What is the trig identity for cos^2x=(half angle)??? help!

OpenStudy (anonymous):

\[\cos(2x)\] or \[\cos^2(x)\]?

OpenStudy (anonymous):

@satellite73 ???

OpenStudy (anonymous):

which one is it?

OpenStudy (anonymous):

oh, it's the latter

OpenStudy (anonymous):

the only thing i know to do with \(\cos^2(x)\) is to rewrite is as \(1-\sin^2(x)\)

OpenStudy (abb0t):

O.O!

OpenStudy (anonymous):

I saw the formula, but I don't know how you would derive that... and especially when you also have a double angle thrown in there... LOL

OpenStudy (anonymous):

it has nothing to do with double angle it comes from the basic \[\sin^2(x)+\cos^2(x)=1\]

OpenStudy (anonymous):

the half angle formula comes from that?!? relly?

OpenStudy (anonymous):

how

OpenStudy (anonymous):

i didn't mention the half angle formula

OpenStudy (anonymous):

the half angle formula comes from taking the double angle formula for cosine and replacing \(2x\) by \(x\)

OpenStudy (anonymous):

...

OpenStudy (anonymous):

?

OpenStudy (anonymous):

You mean the cos2x=cos^2 x - sin^2 x ?

OpenStudy (anonymous):

\[\cos(2x)=\cos^2(x)-\sin^2(x)\] replace \(2x\) by \(x\) and \(x\) therefore by \(\frac{x}{2}\)

OpenStudy (anonymous):

then are you supposed to square it?

OpenStudy (anonymous):

Ok I plugged them in... Now what?

OpenStudy (anonymous):

ok hold the phone i gave you the wrong formula to use

OpenStudy (anonymous):

\[\cos(2x)=2\cos^2(x)-1\] use that one

OpenStudy (anonymous):

oh, ok. Will the other one not work even though they're the same?

OpenStudy (anonymous):

no the other has both sine and cosine no help

OpenStudy (anonymous):

oh, ok that makes sense LOL

OpenStudy (anonymous):

you get \[\cos(x)=2\cos^2(\frac{x}{2})-1\] and you solve this for \(\cos(\frac{x}{2})\)

OpenStudy (anonymous):

ok done

OpenStudy (anonymous):

ahhh and THIS is where you need \(\pm\)!!

OpenStudy (anonymous):

Oh! hahaha good thing I asked that before hand

OpenStudy (anonymous):

wait... where does the plus and minus fit in?

OpenStudy (anonymous):

I got cos^2 (x/2)=[ cos(x)+1 ]/2

OpenStudy (anonymous):

yeah right

OpenStudy (anonymous):

oh... the next step

OpenStudy (anonymous):

but you want to get rid of the square yes next step

OpenStudy (anonymous):

but be careful even though there is a \(\pm\) in the formula, doesn't means you get two answers you have to decide if it is plus or minus

OpenStudy (anonymous):

by looking at the domain that they give you in the problem, right? What do you look at to determine that?

OpenStudy (anonymous):

This is what I got: cos(x/2)= (+-) ROOT[(cosx +1)/2]

OpenStudy (anonymous):

what quadrant your angle is in

OpenStudy (anonymous):

\[\cos(x)=\pm\sqrt{\frac{\cos(x)+1}{2}}\]

OpenStudy (anonymous):

course it is on the cheat sheet as well

OpenStudy (anonymous):

actually it is really good to know how to derive these because who can memorize all this drek easy to know a few and then get the ones you need

OpenStudy (anonymous):

yes, but why does that one have a double angle in it? even if you take the root, there's a double angle in it...

OpenStudy (anonymous):

you lost me

OpenStudy (anonymous):

yeah, it's what I do with my sin^2 +cos^2=1

OpenStudy (anonymous):

ok wait a sec

OpenStudy (anonymous):

cos^2 x=(1/2) (1+cos(2x))

OpenStudy (anonymous):

thats wat it says on the sheet

OpenStudy (anonymous):

fine then take the square root and get exactly what you got before

OpenStudy (anonymous):

oh, and replace \(x\) by \(\frac{x}{2}\)

OpenStudy (anonymous):

oh! So THAT's what happened! LOL that was simple.

OpenStudy (anonymous):

I was like, why is there a double angle in this formula..? hahaha

OpenStudy (anonymous):

yeah simple enough

OpenStudy (anonymous):

why trig so early in the semester?

OpenStudy (anonymous):

Oh, and I need some geomeetry brushing up to do, I'll open a new question for it... I can memorize these geo. volume and are formulas all I want, but I won't understand thme and be able to use them...

OpenStudy (anonymous):

have fun

OpenStudy (anonymous):

Are you leaving?

OpenStudy (anonymous):

:) thanks for your time~ I really appreciate it a lot!